### v2

@@ 4.3 Goodness of Fit Evaluation
In this manuscript, we adapt a third approach, originally proposed in the survival literature. Lin et al. (1993); Borgan and Zhang (2015) introduce a temporal process consisting of cumulative sums of martingale residuals for the Cox proportional hazards model. Under the null hypothesis, this sum has a known distributional behaviour. Given that the presence of time-varying effects in our model formulation violates the proportional hazard assumption, our proposal further relies on the goodness-of-fit tools presented in Marzec and Marzec (1997). Specifically, we rely on a zero-mean Martingale-residual type process $G[\hat{\bm{\gamma}},\cdot|\mathcal{E}]$, defined as a weighted cumulative sum of Martingale residuals $\widehat{M}_{sr}=N_{sr}(t)-\widehat{\Lambda}_{sr}(t)$, for any statistic of interest $\phi$, and evaluated at $n$ equally spaced time points $u\in[0,1]$:
@@
The curve $G$ measures the difference between the observed statistic $\phi$ and its expected value across time. In particular, we are interested in testing whether the covariates are included appropriately in the model, i.e., $\phi=x^{j}$. In this case, $G[\hat{\bm{\gamma}},\cdot|\mathcal{E}]$ can be shown to be a bridge process with $G[\hat{\bm{\gamma}},0|\mathcal{E}]=G[\hat{\bm{\gamma}},1|\mathcal{E}]=0$.
Given an estimate for the variance of an individual contribution to process $G[\hat{\bm{\gamma}},\cdot|\mathcal{E}]$, named $\hat{J}_{G[\hat{\bm{\gamma}}]}$, it is possible to implement a formal, Kolmogorov-Smirnov (KS) type statistical test:
@@
$\hat{\bm{J}}_{G}=n^{-1}\times\sum_{k=1}^{n}\bm{G}^{\bm{i}}_{s_{k}r_{k}}[\hat{\bm{\gamma}},t_{k}]\bm{G}^{\bm{i}}_{s_{k}r_{k}}[\hat{\bm{\gamma}},t_{k}]^{T}$
If $\bm{G}^{\bm{i}}[\hat{\bm{\gamma}},\cdot|\mathcal{E}]$ is the unpenalised score vector, one can use instead the observed Fisher information matrix.
It follows that the scaled score vector,
@@
converges to a a multivariate Brownian bridge $\bm{Z}^{0}$ under the assumption that the model is correctly incorporating the covariate (Hjort and Koning, 2002).
To test the fit of a covariate whose elements in the model matrix correspond to $\bm{h}_{\bm{i},sr}$, we propose the following test-statistic,
@@
Refer to caption Refer to caption
@@
Figure 1: a) Comparison between the true and the estimated coefficients on the simulated data; b) Comparison between the true (undashed line) and the non-parametric estimates (dashed) of the cumulative baseline hazard. True baseline hazard is assumed to be constant and equal to $0.008$. c) Comparison between the true and estimated random effects for species invasiveness and region invasibility. True values are represented by the conditional expectation of random effects fit on real data, while estimates correspond to the $0-$ dimensional-spline-estimates on the simulated data. Comparison has been performed in terms of empirical cumulative distribution function. d) Empirical distribution of the p-values resulting from testing climatic dissimilarity, distance and their global test. The empirical distributions (yellow, blue, and brown respectively) are compared with the uniform cumulative distribution (black), which is the expected distribution of the p-values when the model is adequate.
@@ 5 Simulation Study
In order to show that the available data and model formulation are adequate to draw meaningful conclusions, we provide a brief simulation study to evaluate (i) the estimation of the mixed additive relational event model, (ii) the non-parametric technique for evaluating the baseline hazard function, and (iii) the strategy for assessing the goodness-of-fit within a data-setting similar to the one used in practice. To achieve this, we first fit a basic model that includes distance, climate dissimilarity for insect invasions, obtaining estimates for the time-varying effect of the former, and a constant effect of the latter. We also include random species and regions effects. We define a native range $\mathcal{NR}(1880)$ similar to the observed process and set the risk set in 1880 as its complement, $\mathcal{R}(1880)=\mathcal{NR}(1880)^{c}$.
We then proceed to simulate complete IPs $210$ times for all insect species, using these parameter and random effect values over a time period from 1880 to 2005 via Gillespie-type algorithm (Gillespie, 1977). Inter-arrival times are simulated with piecewise constant rates,
@@
For each simulated data set, we perform a NCC inference procedure, estimating the fixed parameters and random effects, the baseline hazard, and performing goodness-of-fit estimates. We sample a case-control data set, with the number of rows equal to the number of simulated events, a response column of $1$s and columns with covariate differences $\Delta\bm{x}$ and $\Delta\bm{z}$ for the event and randomly sampled non-event. We fit a additive mixed-effect logistic regression to obtain the parameter estimates $\hat{\beta}_{dt},\hat{\beta}_{d}(t)$ and conditional expectations of the random effects $\hat{b}_{s},\hat{b}_{r}$. Figure 1 a) and c) show that NCC estimation procedure achieves unbiased results. Estimates for cumulative baseline hazard in Figure 1 b) show that the method is adequately capturing the underlying trend. Finally, the goodness-of-fit plots in Figure 1 for climatic similarity, distance, and their global test illustrate that the proposed method adequately captures the model’s goodness-of-fit.
@@ 6 Analysis of Plant and Insect Co-invasions
The main substantial goal of this paper is to sharpen our understanding of which forces shape the diffusion of alien species across the globe. This is a complex process with many feedback loops that have previously been ignored. By properly accounting for the endogenous and temporal nature of the process, our model captures the complexity of the joint insect and plant invasion process. Plants and insects have diverse and complex relationships that can be beneficial, harmful, or neutral.
Beneficial interactions between plants and insects are typically referred to as mutualism (Bronstein et al., 2006). Many plants rely on insects for pollination, where insects transfer pollen from the male reproductive organs (stamens) of one flower to the female reproductive organs (pistils) of another flower. Insects such as bees, butterflies, moths, flies, and beetles are important pollinators. They obtain nectar or pollen as a food source from flowers while inadvertently transferring pollen. Some insects help in the seed dispersal of plants. They may consume fruits or seeds and then transport them to different locations, aiding in the plant’s colonization and distribution. Examples include ants that carry seeds to their nests and bury them. In a process called myrmecophytes, plants provide shelter and food resources such as nectar or specialized structures called extrafloral nectaries, while the ants, ladybugs and wasps offer protection from herbivores and competing plants (Project Learning Tree, 2023). Insects can aid in nutrient cycling by breaking down plant material or dead organisms, promoting decomposition and recycling of organic matter. For instance, certain beetles and flies help break down decaying plant matter, contributing to soil fertility (Uma et al., 2023).
@@
Not all interactions between plants and insects are mutually beneficial. Some insects, like herbivores or plant pathogens, can harm plants by feeding on their tissues or spreading diseases. In response, plants may develop defense mechanisms such as producing toxins, thorns, or chemical signals to repel or deter these insects (War et al., 2012). Many insects are herbivores and feed on plant parts such as leaves, stems, flowers, or fruits. Herbivory can cause direct physical damage to plants, leading to reduced growth, impaired photosynthesis, or even death. Some insects, such as gypsy moth larvae and tent caterpillars, specialize in defoliating plants by consuming their leaves (Hemming and Lindroth, 1995), whereas other insects, such as aphids and mealybugs, feed on plant sap by inserting their mouthparts into plant tissues and extracting nutrients (Branco et al., 2023). Some insects consume plant seeds, affecting the plant’s reproductive success. Another type of non-mutualistic interaction is gall formation. Certain insects induce the formation of abnormal growths, called galls, on plants, manipulating its physiology to create a protective structure that provides the insect with food and shelter (Takeda et al., 2021). Finally, insects can act as vectors of disease transmission. They may pick up pathogens from infected plants, and transmit them to healthy plants while feeding or through physical contact. Aphids play a central role in virus transmission, that may lead to plant diseases, manifesting, for example, with yellowing (Sankarganesh et al., 2020).
@@
Figure 2: Values of AIC for the examined model formulations. We outline that, whenever included, distance, trade, and agricultural land-coverage are supposed to have a time-varying impact while climatic dissimilarity, urban land-coverage, and colonial ties a fixed effect. This choice come from the previous studies on the topic (Juozaitienė et al., 2023). On the other side, all the considered models include random effects for species invasiveness, region invasibility and species co-invasion. The best model in terms of corrected AIC (Wood et al., 2016) includes distance, trade, colonial ties and climatic dissimilarity and is outlined in the plot with a red crossed symbol. According to the covariates included in the compared models, we can distinguish seven groups of model formulations.
Differential mutualisms between plants and insects seem to have a role in alien species invasions (Prior et al., 2015; Simberloff and Von Holle, 1999). The term invasional meltdown has been introduced for how non-indigenous species may collaborate to increase the likelihood of a successful invasion. For instance, non-native plants and insects may cause variations in their new habitats enhancing the opportunity for other non-indigeneous species to get established in these areas.
@@
One aim is to study alien plant and insect co-invasions. In addition to the drivers reported in Table 1, we introduce the three random effects. Besides the two main effects, species invasiveness and region popularity, we define a species interaction effect, $b_{ss^{\prime}}$, which captures how much the presence of species $s^{\prime}$ in a certain region stimulates or impedes the invasion by species $s$ into that region. These two entries can assume the categories Rare interaction and Novelty as well. The former is used when the species-last species interaction appears only once among those recorded for events and non-events; the latter is taken into account if no other species are detected in the related country before the considered time. We point out that species and last species, in this context, may be members of separate taxonomies, allowing us to study possible symbiotic relationships between insects and vascular plants.
Model selection has been conducted by evaluating the corrected version of AIC (Wood et al., 2016). This correction, relying on the adjustment for the degrees of freedom, allows avoiding the tendency to chiefly select the simplest or the more complex model. Figure 2 shows the values of AIC for the $63$ evaluated model formulations, including different subsets of the covariates shown in Table 1. Following Juozaitienė et al. (2023), whenever included, distance, trade and agricultural land-coverage have a time-varying effect, while climatic dissimilarity, urban land-coverage and colonial ties are supposed to remain constant over time. The best model according to corrected AIC includes distance, trade, colonial ties and difference in temperature.
@@
Land-cover characteristics seem to have no impact on the dynamics of invasions. Both the fixed coefficients, related to the climatic conditions and to the spread of the colonialism turn out to be negative. The former result is quite intuitive: species tend to invade regions that have similar climatic characteristics as the regions they have invaded before. The latter instead is somewhat counter-intuitive, but it could be a residual effect as trade has a strong positive effect. However, as shown in Figure 3b, the effect of trade seems to be diminishing over the last century, surprisingly. Juozaitienė et al. (2023) suggest that it is now more focused on products than on raw materials. Additionally, the augmented control at the borders may have led to a decrease of the effect trade on alien invasions. The time-varying effect related to distance is negative as seen in Figure 3a, and relatively constant between 1880 and 2005. This fact confirms how unusual long-distance natural invasion occurrences are.
Refer to caption Refer to caption
Figure 3: Time-varying coefficients for distance (Left) and trade (Right) (undashed lines) with the related posterior confidence intervals (dashed lines).
Once the fixed and the time-varying effect of the covariates are properly accounted for, we may look to the largest estimated values of the random effects related to the insect and plant invasiveness. Frankliniella occidentalis, the most invasive insect found, has a long history of foreign species incursions dating back to the $1970$s. It is widely known for its pest qualities and ability to cause significant plant harm (Kirk and Terry, 2003). Another largely invasive species is Anoplolepis gracilipes, a very small ant, which is able to negatively impact on the native ecosystems, particularly on forests. On the other hand, among the most invasive plants represented in Figure 5, we can identify Chromolaena odorata. While its effect in South Africa have been largely studied and need to be controlled, it also has beneficial effect in central Africa. Insect and plant invasiveness random effects have, respectively, standard deviations of $1.86$ and $0.31$, suggesting that the heterogeneity on the insect side is markedly higher than on the plant side. Figure 4 depicts the regions’ invasibility by means of their conditional random effects. Australia and Canada have the highest invasibility, but also South Africa, United States, and New Zealand have high values, meaning that the probability of alien species invasions in these regions is larger than in other areas.
Figure 4: Regions’ invasibility in terms of estimated random effects. Lighest-blue areas are those that the model identifies as most popular, such as Australia and Canada. On the other hand, darker locations are those that lead to a decrease in the rate of occurrence of alien species invasions (Peru and Saudi Arabia are some instances).
Figure 5 illustrates the strongest co-invasion relationships between the species. Particularly, it shows whether the presence of a species affects the likelihood of an invasion event by another species. It is intriguing to consider the interactions between species that pertain to different taxa: for instance, we can see that a positive effect is estimated for Phenacoccus manihoti, commonly known as the cassava mealybug, when Chromolaena odorata or Siam weed has reached the country. Calatayud et al. (1994) studied the variations in the dispersion dyamics of P. manihoti in relation to other factors in Brazzaville (Congo), where C. odorata is the primary plant species. The opposite tendency, i.e., a negative interaction is estimated for Frankliniella occidentalis (western flower thrip) invading a region that has already been reached by Achyranthes aspera (chaff-flower). We can find in the literature two instances where the aforementioned weed seem related to viruses whose transmission F. occidentalis may be involved. Particularly, A. aspera is specifically mentioned by Kumar et al. (2008) as being a part of the Tobacco Streak Virus’ native range. Because of the harm caused by this virus, one potential response may be the development of a resistance, which has actually been observed in a plant that is resistant to F. occidentalis. Additionally, Macharia et al. (2016) analysis of the plants in the tomato production area in Kenya revealed the presence of A. aspera. However, it is unclear how it might act as a host for the Tomato spotted wilt virus, for which F. occidentalis is a renowned vector.
Once all the features have been taken into account, we can inspect the estimated cumulative baseline hazard. In Figure 6, one can find the estimates according to Equation (8). The shape of the curve for insects suggests that the rate of invasions for this taxonomy is growing in a way that the model’s built-in drivers are unable to account for. Yet, for vascular plants, the existence of a linear estimated function may be interpreted as a constant underlying baseline. This result confirms the literature (Bonnamour et al., 2021). They outline the largest increase in the rate for insects comparing to plants. Particularly, they claim that different scale is due to the fact that insects took more advantage from the availability of fast transportation tools rather than plants, already able to survive for longer time across journeys. Futhermore, due to the smaller dimension of insects and to their mobility, insects may also be harder to identify (Walliser, 2013).
Given our selected model, we now aim to understand if the included covariates correctly incorporate and describe the dynamics driving the alien species invasions. For this we will use the goodness-of-fit technique presented before. Particularly, when dealing with Colonial Ties and Climatic Dissimilarity, KS test in Equation 10 can be applied. Instead, for Distance and Trade, we employ the KS test in Equation (11), with $q=10$. We remark that the effects of these two covariates are assumed to be $10$-dimensional non-linear functions of time, so $10$ components of the score will refer to each of these two covariates. Table 3 reports the p-values of the associated tests.
As we can see from Table 3, the global test in Equation (12) is adequate, with a $p$-value equal to $0.468$. In principle, this means that overall the model is adequate and no further tests have to be performed. If, for argument’s sake, we would check the individual tests, we see that all covariates except distance are adequately incorporated in the model. When the distance test leads to rejection, we can look at the graphical representation of the score process. As mentioned above, the $p$-value is obtained empirically, by emulating the theoretical behaviour of the multivariate Brownian Bridge. We can thus compare graphically the observed $\lVert\bm{M}[\hat{\bm{\gamma}},u|\mathcal{E}]\rVert^{2},u\in[0,1]$ against the squared norms of Brownian Bridges $\sup_{u\in[0,1]}{\lVert\bm{Z}^{0}(u)\rVert^{2}}$, as shown in Figure 6b.
@@
Figure 6: a) Matched and Pooled Estimates of the Cumulative Baseline Hazard. Following Borgan and Langholz (1997), we may provide a taxa-specific estimate of the baseline (blue for insects, red for plants); as an alternative, we may rely on a pooled estimate (yellow line) that takes into account the fact that the sampled control belongs to the same group as the observed case. b) Goodness of Fit of Distance Covariate. Graphical comparison between the observed $\lVert\bm{M}[\hat{\bm{\gamma}},u|\mathcal{E}]\rVert^{2},u\in[0,1]$ against $200$ simulated squared norm of Brownian Bridges $\sup_{u\in[0,1]}{\lVert\bm{Z}^{0}(u)\rVert^{2}}$. We plot the time in the original scale in order to understand which section of the time-window may lead to misspecification. It is interesting to see that the Bridge overcome the cloud created by the theoretical curves in the time range between $1961$ and $1983$ and in the time window between $1993$ and $2004$.

### v3

@@ 4.3 Goodness of Fit Evaluation
In this manuscript, we adapt a third approach, originally proposed in the survival literature. Lin et al. (1993); Borgan and Zhang (2015) introduce a temporal process consisting of cumulative sums of martingale residuals for the Cox proportional hazards model. Under the null hypothesis, this sum has a known asymptotic distributional behaviour. However, given that the presence of time-varying effects in our model formulation violates the proportional hazard assumption, our proposal further relies on the goodness-of-fit tools presented in Marzec and Marzec (1997). Specifically, we rely on a zero-mean Martingale-residual type process $G[\hat{\bm{\gamma}},\cdot|\mathcal{E}]$, defined as a weighted cumulative sum of Martingale residuals $\widehat{M}_{sr}=N_{sr}(t)-\widehat{\Lambda}_{sr}(t)$, for any statistic of interest $\phi$, and evaluated at $n$ equally spaced points $u\in[0,1]$:
@@
The curve $G$ measures the difference between the observed statistic $\phi$ and its expected value across time. In particular, we are interested in testing whether the covariates are included appropriately in the model, i.e., $\phi=x^{j}$. In this case, $G[\hat{\bm{\gamma}},\cdot|\mathcal{E}]$ can be shown to be a bridge process with $G[\hat{\bm{\gamma}},0|\mathcal{E}]=G[\hat{\bm{\gamma}},1|\mathcal{E}]=0$. Given an estimate for the variance of an individual contribution to process $G[\hat{\bm{\gamma}},\cdot|\mathcal{E}]$, named $\hat{J}_{G[\hat{\bm{\gamma}}]}$, it is possible to implement a formal, Kolmogorov-Smirnov (KS) type statistical test:
@@
$\hat{\bm{J}}_{G}=n^{-1}\times\sum_{k=1}^{n}\bm{G}^{\bm{i}}_{s_{k}r_{k}}[\hat{\bm{\gamma}},t_{k}]\bm{G}^{\bm{i}}_{s_{k}r_{k}}[\hat{\bm{\gamma}},t_{k}]^{\prime}$
If $\bm{G}^{\bm{i}}[\hat{\bm{\gamma}},\cdot|\mathcal{E}]$ is the unpenalised score vector, one can use instead the observed Fisher information matrix. It follows that the scaled score vector,
@@
converges to a a multivariate Brownian bridge $\bm{Z}^{0}$ under the assumption that the model is correctly incorporating the covariate (Hjort and Koning, 2002). To test the fit of a covariate whose elements in the model matrix correspond to $\bm{h}_{\bm{i},sr}$, we propose the following test-statistic,
@@
Refer to caption Refer to caption Refer to caption
@@
Figure 1: Simulation Study: Summary of the Results. a) Comparison between the true coefficients (black solid) and the estimated coefficients for climatic dissimilarity (yellow dashed) and distance (blue dashed) on the simulated data. b) Comparison between the true (black solid) and the non-parametric estimates (blue dashed) of the cumulative baseline hazard. True baseline hazard is assumed to be constant and equal to $0.008$. c) Comparison between the true (black solid) and estimated random effects for species invasiveness (yellow) and region invasibility (blue). True values are represented by the conditional expectation of random effects fit on real data, while estimates correspond to the 00-dimensional spline estimates on the simulated data. Comparison has been performed in terms of empirical cumulative distribution function. d) Empirical distribution of the p-values resulting from testing climatic dissimilarity (yellow), distance (blue) and their global test (brown). The empirical distributions are compared with the uniform cumulative distribution (black), which is the expected distribution of the p-values when the model is adequate.
@@ 5 Simulation Study
In order to show that the available data and model formulation are adequate to draw meaningful conclusions, we provide two simulation studies. First we would like to evaluate the inference techniques proposed in this paper for the type of empirical setting considered. Secondly, we want to evaluate the possible effect of model misspecification. As the recording of invasive species may not be uniform across the globe, we want to evaluate this potential recording bias on the parameter estimates in the model.
5.1 Validation of Introduced Techniques
In this section, we aim to evaluate (i) the estimation of the mixed additive relational event model, (ii) the non-parametric technique for evaluating the baseline hazard function, and (iii) the strategy for assessing the goodness-of-fit within a data-setting similar to the one used in practice. To achieve this, we first fit a basic model that includes distance, climate dissimilarity for insect invasions, obtaining estimates for the time-varying effect of the former, and a constant effect of the latter. We also include random species and regions effects. We define a native range $\mathcal{NR}(1880)$ similar to the observed process and set the risk set in 1880 as its complement, $\mathcal{R}(1880)=\mathcal{NR}(1880)^{c}$.
We then proceed to simulate complete IPs $210$ times for all insect species, using these parameter and random effect values over a time period from 1880 to 2005 via Gillespie-type algorithm (Gillespie, 1977). Inter-arrival times are simulated from a continuous exponential distribution with piecewise-constant rates,
@@ 5.1 Validation of Introduced Techniques
For each simulated data set, we perform a NCC inference procedure, estimating the fixed parameters and random effects, the baseline hazard, and performing goodness-of-fit estimates. We sample a case-control data set, with the number of rows equal to the number of simulated events, a response column of $1$s and columns with covariate differences $\Delta\bm{x}$ and $\Delta\bm{z}$ for the event and randomly sampled non-event. Furthermore, time is discretised in a way simulated data can match actual data as closely as possible. We fit a additive mixed-effect logistic regression to obtain the parameter estimates $\hat{\beta}_{dt},\hat{\beta}_{d}(t)$ and conditional expectations of the random effects $\hat{b}_{s},\hat{b}_{r}$. Figure 1 a) and c) show that NCC estimation procedure achieves unbiased results. Estimates for cumulative baseline hazard in Figure 1 b) show that the method is adequately capturing the underlying trend. Finally, the goodness-of-fit plots in Figure 1 for climatic similarity, distance, and their global test illustrate that the proposed method adequately captures the model’s goodness-of-fit. There is no evidence that the discretisation of time is negatively impacting the quality of the results.
5.2 Recording Bias
The second simulation study evaluates the behaviour of our estimation procedure in the presence of recording bias. As highlighted by Bonnamour et al. (2021), variation in scientific recording efforts can influence our understanding of alien species diffusion dynamics. To assess this effect, we simulate species invasion as a function of an endogenous driver, e.g., ecological similarity between the destination region and the most recent source region where the species was detected. To account for differences in sampling effort across regions, we introduce a masking probability for each region and randomly remove a portion of the simulated invasions accordingly. Figure 2 presents the estimated coefficients from 100 experimental replications against the true parameter used for data generation. Even under strong masking probabilities, our methodology correctly identifies the sign of the coefficients. In the empirical application, we mainly focus on the interpretation of the sign and trend of the effects, rather than their absolute magnitudes. We also observe that increasing masking strength shrinks coefficient estimates toward zero. Nevertheless, even when the majority of the records from certain regions is missing, the direction of the effect remains correctly identified and does not overlap with zero for any instance.
Figure 2: Simulation Study on Recording Bias. Estimated coefficients from 100 experiment replications are compared to the true parameters used in data simulation. The maximum masking probability represents the upper limit of region-specific masking probabilities, ranging from 0 to this maximum value. While increasing the maximum masking probability shrinks the estimates towards zero, the sign of the effect remains correctly identified.
@@ 6 Analysis of Plant and Insect Co-invasions
The main substantial goal of this paper is to understand which forces shape the diffusion of alien species across the globe. This is a complex process with many feedback loops that have typically been ignored in previous studies. By properly accounting for the endogenous and temporal nature of the process, our model captures the complexity of the joint insect and plant invasion process. Plants and insects have diverse and complex relationships.
@@
Beneficial interactions between plants and insects are typically referred to as mutualism (Bronstein et al., 2006). However, not all interactions between plants and insects are mutually beneficial. Many insects are herbivores and feed on plant parts such as leaves, stems, flowers, or fruits. Herbivory can cause direct physical damage to plants, leading to reduced growth, impaired photosynthesis, or even death. Some insects, such as spongy moth larvae and tent caterpillars, specialize in defoliating plants by consuming their leaves (Hemming and Lindroth, 1995), whereas other insects, such as aphids and mealybugs, feed on plant sap by inserting their mouthparts into plant tissues and extracting nutrients (Branco et al., 2023). Some insects consume plant seeds, affecting the plant’s reproductive success. Another type of non-mutualistic interaction is gall formation. Certain insects induce the formation of abnormal growths, called galls, on plants, manipulating its physiology to create a protective structure that provides the insect with food and shelter (Takeda et al., 2021). Finally, insects can act as vectors of disease transmission. They may pick up pathogens from infected plants, and transmit them to healthy plants while feeding or through physical contact. Aphids play a central role in virus transmission, that may lead to plant diseases, manifesting, for example, with yellowing (Sankarganesh et al., 2020). Differential mutualisms between plants and insects seem to have a role in alien species invasions (Prior et al., 2015; Simberloff and Von Holle, 1999). The term invasional meltdown has been introduced for how non-indigenous species may collaborate to increase the likelihood of a successful invasion. For instance, non-native plants and insects may cause variations in their new habitats enhancing the opportunity for other non-indigeneous species to get established in these areas.
@@
Figure 3: Model Selection. Values of AIC for the examined model formulations. We outline that, whenever included, distance, trade, and agricultural land-coverage are supposed to have a time-varying impact while climatic dissimilarity, urban land-coverage, and colonial ties a fixed effect. This choice comes from the previous studies on the topic (Juozaitienė et al., 2023). On the other side, all the considered models include random effects for species invasiveness, region invasibility and species co-invasion. The best model in terms of corrected AIC (Wood et al., 2016) includes distance, trade, colonial ties and climatic dissimilarity and is outlined in the plot with a red crossed symbol. According to the covariates included in the compared models, we can distinguish seven groups of model formulations.
@@
One of the aims is to study alien plant and insect co-invasions. In addition to the drivers reported in Table 1, we introduce the three random effects. Besides the two main effects, species invasiveness and region popularity, we define a species interaction effect, $b_{ss^{\prime}}$, which captures how much the presence of last species $s^{\prime}$ in a certain region stimulates or impedes the invasion by species $s$ into that region. These two entries can assume the categories Rare interaction and Novelty as well. The former is used when the species-last species interaction appears only once among those recorded for events and non-events; the latter is taken into account if no other species are detected in the related country before the considered time. We point out that species and last species, in this context, may be members of separate taxonomies, allowing us to study possible symbiotic relationships between insects and vascular plants. The species interaction effect further underlines that the conditional independence assumption does not imply independence between events in this model, as past events can influence current ones. However, these relationships are assumed to be fully explainable through the history of previous occurrences.
@@
Model selection has been conducted by evaluating the corrected version of AIC. This correction, relying on the adjustment for the degrees of freedom, avoids chiefly selecting the simplest or the most complex model. Figure 3 shows the values of AIC for the $63$ evaluated model formulations, including different subsets of the covariates shown in Table 1. Following Juozaitienė et al. (2023), whenever included, distance, trade and agricultural land-coverage have a time-varying effect, while climatic dissimilarity, urban land-coverage and colonial ties remain constant over time. The best model according to corrected AIC includes distance, trade, colonial ties and difference in temperature. Land-cover characteristics seem to have no impact on the dynamics of invasions. Both the fixed coefficients related to the climatic conditions and to the spread of the colonialism turn out to be negative. The former result is quite intuitive: species tend to invade regions that have similar climatic characteristics as the regions they have invaded before. The latter instead is somewhat counter-intuitive, but it could be a residual effect as trade has a strong positive effect. Perhaps surprisingly, as shown in Figure 4b, the effect of trade seems to be diminishing over the last century. Juozaitienė et al. (2023) suggest that this may be the result from the fact that international trade is now more focused on products rather than on raw materials. Additionally, stricter international transportation regulations may have led to a decrease of the effect of trade on species invasions. The time-varying effect related to distance is negative as seen in Figure 4a, and remains relatively constant between 1880 and 2005. This fact confirms how unusual long-distance invasion occurrences are.
@@
Figure 4: Time-varying estimates. Time-varying estimated coefficients for a) distance and b) trade (undashed lines) with the related posterior confidence intervals (dashed lines).
Once the fixed and the time-varying effects of the covariates are properly accounted for, one can consider the conditional expectation of the random effects related to the insect and plant invasiveness. Frankliniella occidentalis, the most invasive insect found, has a long history of foreign species incursions dating back to the $1970$s. It is widely known for its pest qualities and ability to cause significant plant harm (Kirk and Terry, 2003). Another largely invasive species is Anoplolepis gracilipes, a very small ant, which negatively impacts native ecosystems, particularly forests (Lee and Yang, 2022). Among the most invasive plants represented in Figure 6 is Chromolaena odorata. While its negative effects in South Africa have been studied extensively, leading to its classification as a controlled species, it has also been noted for its beneficial effects in central Africa (Goodall and Erasmus, 1996). The random effects of insect and plant invasiveness have, respectively, standard deviations of 1.86 and 0.31, suggesting that the heterogeneity on the insect side is markedly higher than on the plant side. Figure 5 depicts the regions’ invasibility by means of their conditional random effects. Australia and Canada have the highest invasibility, but also South Africa, United States, and New Zealand have high values, meaning that the probability of alien species invasions in these regions is larger than in other areas.
Figure 5: Regions’ invasibility in terms of estimated random effects. Lightest-blue areas are those that the model identifies as most popular, such as Australia and Canada. On the other hand, darker locations are those that lead to a decrease in the rate of occurrence of alien species invasions (Peru and Saudi Arabia are some instances).
Figure 6 also illustrates the strongest co-invasion relationships between the species. Particularly, it shows whether the presence of a species affects the rate of an invasion event by another species. Considering interactions between species that pertain to different taxa, we see a positive effect for Phenacoccus manihoti, commonly known as the cassava mealybug, when Chromolaena odorata, or Siam weed, has reached the country. Calatayud et al. (1994) studied the variations in the dispersion dyamics of P. manihoti in relation to other factors in Brazzaville (Congo), where C. odorata is the primary plant species. The opposite tendency, i.e., a negative interaction effect, is found for Frankliniella occidentalis (western flower thrip) invading a region that has already been reached by Achyranthes aspera (chaff-flower). We can find in the literature two instances where the aforementioned weed seem related to viruses whose transmission F. occidentalis may be involved. Particularly, A. aspera is specifically mentioned by Kumar et al. (2008) as being a part of the Tobacco Streak Virus’ native range. Because of the harm caused by this virus, one potential response may be the development of a resistance, which has actually been observed in a plant that is resistant to F. occidentalis. Additionally, the analysis of the plants in the tomato production area in Kenya in Macharia et al. (2016) revealed the presence of A. aspera. However, it is unclear how it might act as a host for the Tomato spotted wilt virus, for which F. occidentalis is a renowned vector.
Once all the features have been taken into account, Figure 7 represents the estimated cumulative baseline hazard according to (8). The shape of the curve for insects suggests that the rate of invasions for this taxonomy has been growing strongly in recent years, in a way that the above covariates are unable to account for. Yet, for vascular plants, the existence of a roughly linear baseline hazard may be interpreted that most of the important covariates for plants are accounted for in the selected model. This result is in line with similar results reported in the literature. Bonnamour et al. (2021) report that the invasive behaviour of insects in the past decades has been stronger than that of plants. Particularly, they claim this is due to the fact that insects took more advantage from the availability of fast transportation tools rather than plants. Moreover, insects are able to survive for longer across long journeys. Walliser (2013), furthermore, argued that due to their smaller dimensions and inherent mobility insects may also be harder to spot and control than plants.
Covariate p-value Dimension $q$
distance 0.0002 10
trade 0.580 10
colonial ties 0.404 1
climatic dissimilarity 0.480 1
Omnibus test 0.468
Finally, we aim to understand if the included covariates correctly incorporate and describe the dynamics driving the alien species invasions. For this we will use the goodness-of-fit technique presented above. We use the KS test in (10) to test the fixed linear effects Colonial Ties and Climatic Dissimilarity. Instead, for Distance and Trade, we employ the KS test in (11) with $q=10$, as these effects are modelled via $10$-dimensional non-linear functions of time. Table 3 reports the p-values of the associated tests. Importantly, the global test (12) is not rejected, with a $p$-value equal to $0.468$. In principle, this means that overall the model is adequate and no further tests have to be performed. If, for argument’s sake, we would check the individual tests, we see that all covariates except distance are adequately incorporated in the model. We hypothesize that the effect of distance might have changed in more recent times, as a result of faster and more varied modes of transport. Although we allow the effect of distance to change over time, i.e., $\beta_{d}(t)\cdot d_{sr}(t)$, this may not have been sufficient and that the effect of distance change both over time and over distance itself in a non-linear way, i.e., $f\left(t,d_{sr}(t)\right)$. As mentioned above, the $p$-value is obtained empirically, by emulating the theoretical behaviour of the multivariate Brownian Bridge. We can thus compare graphically the observed $\lVert\bm{M}[\hat{\bm{\gamma}},u|\mathcal{E}]\rVert^{2},u\in[0,1]$ against the squared norms of Brownian Bridges $\sup_{u\in[0,1]}{\lVert\bm{Z}^{0}(u)\rVert^{2}}$, as shown in Figure 7b.
Refer to caption Refer to caption
(a) (b)
@@
Figure 7: a) Matched and Pooled Estimates of the Cumulative Baseline Hazard. Following Borgan and Langholz (1997), we may provide a taxa-specific estimate of the baseline (blue for insects, yellow for plants, solid); as an alternative, we may rely on a pooled estimate (dotted) that takes into account the fact that the sampled control belongs to the same group as the observed case. b) Goodness of Fit of Distance Covariate. Graphical comparison between the observed squared norm of the martingale-residual process against $200$ simulated squared norm of Brownian Bridges. We plot the time in the original scale in order to understand which section of the time-window may lead to misspecification. It is interesting to see that the Bridge overcome the cloud created by the theoretical curves in the time range between $1961$ and $1983$ and in the time window between $1993$ and $2004$. We hypothesize that the impact of distance may have evolved in recent times, suggesting that more complex, non-linear, effects of distance, could improve the model fit.